# The Low-Degree Hardness of Finding Large Independent Sets in Sparse Random Hypergraphs

### v2

@@ 3.1.2 From Local to Low-Degree
(Here, $\cong$ denotes the isomorphism of rooted hypergraphs). The following lemma shows that $n_{T}$ is well-concentrated.
@@
The proof follows by combining Markov’s inequality with a hypergraph version of Proposition 12.3 [barak2019nearly] (for which the proof is identical, mutatis mutandis). ∎
@@
The proof follows by a hypergraph version of [barak2019nearly, Lemma 12.4] (for which the proof is identical, mutatis mutandis), and linearity of expectation. ∎
@@
For any $\tau>0$ and $n$ sufficiently large in terms of $d,\,s,\,r,$ and $\tau$, there is a constant $C\coloneqq C(d,s,r,\tau)>0$ such that
@@
Before we prove the main result, we make an analogous observation to the one in [wein2022optimal, Lemma 2.11] on graphs. In particular, a random polynomial can be converted to a deterministic polynomial that works almost as well.
@@
Thus, there exists a $\omega^{*}\in\Omega$ such that the deterministic degree-D polynomial $f^{*}(\cdot)=f(\cdot,\omega)$ satisfies
$\mathbb{E}_{A}[\|f(A)\|^{2}]\leq c\xi k\quad\text{and}\quad\mathbb{P}_{A}[|V_{f}^{\eta}(A)|<k]\leq c\delta$
@@
Given $\varepsilon>0$, let $d$ and $s$ be sufficiently large in order to apply Theorem 3.1. As a result, we obtain an $s$-local function $g$ that outputs independent sets of density
@@
which concludes that the rounding procedure succeeds with probability at least $1-\exp(-\Omega(n^{1/3}))$.
@@
The idea now is to partition $V_{q}$ into subsets $U_{1},\ldots,U_{(r-1)q+1}$ such that for each $U_{i}$, the vertices in $U_{i}$ have disjoint $s$-neighborhoods. In particular, the random variables $\{\mathbbm{1}_{v\in I_{f}}\}_{v\in U_{i}}$ are independent conditioned on the outcome $A$. Each vertex $v\in V_{q}$ has at most $(r-1)q$ vertices in $N_{2s}(A,v)$, which implies there are at most $(r-1)q$ vertices $u\in V_{q}$ such that $u\neq v$ and $N_{s}(A,v)\cap N_{s}(A,u)\neq\varnothing$. We can greedily form the partition $\{U_{i}\}$ while maintaining the disjointness property (given that there are $(r-1)q+1$ many such subsets available). For such a partition $\{U_{i}\}$, we apply Chernoff’s bound for each $i$ to obtain the following:
@@
It suffices to show that $\mathbb{E}_{A,\mathbf{X}}[f_{v}(A,\mathbf{X})^{2}]=O(1)$ for all vertices $v$. Fix a vertex $v$ and let $N=|N_{s}(A,v)|$. By (3), for each $H\in\mathbb{H}_{v,s,q}$ the corresponding term is not 00 only if $H$ is a subgraph of $N_{s}(A,v)$. This implies that number of of nonzero terms is at most
(${N\choose\leq q}=\sum_{i=0}^{q}{N\choose i}\leq\sum_{i=1}^{q}N^{i}\leq(N+1)^{q}.$
@@
In order to bound the $\mathbb{E}_{A,\mathbf{X}}[(f_{v}(A,\mathbf{X})^{2}]$, it is sufficient to prove a tail bound for $N$. Starting from $m_{0}=1$, let $m_{i}$ be the number of vertices whose distance from $v$ in $A$ is exactly $i$. Conditioned on $m_{i}$, $m_{i+1}$ is stochastically dominated by $\mathsf{Bin}\left(r{n-1\choose r-1}m_{i},d/{n-1\choose r-1}\right)$. By a Chernoff bound, for fixed $m_{i}$ and any $\delta\geq 1$, we have
$\mathbb{P}[m_{i+1}\geq(1+\delta)rdm_{i})]\leq\exp\left(-\frac{\delta rdm_{i}}{3}\right)\leq\exp\left(-\frac{\delta rd}{3}\right).$
Taking a union bound over all $i\in[s]$,
@@
For $\delta^{{}^{\prime}}=(3\delta rd)^{s}$, we have an exponentially small tail bound on $N$:
@@ 3.2 Forbidden Structures
In this section, we will define two “forbidden structures” in order to assist with our proof of Theorem 1.6. Before we define these structures and prove the corresponding results, we state the following theorem due to Krivelevich and Sudakov on the size of the maximum independent set in $\mathcal{H}_{r}(n,p)$ (this result wasn’t explicitly stated, but can inferred from Corollary 1 in their paper) [krivelevich1998chromatic].
@@
Let $\varepsilon>0$ and $r\in\mathbb{N}$ such that $r\geq 2$. There is $d^{*}>0$ such that for any $d\geq d^{*}$, there exists $n^{*}>0$ such that for any $n\geq n^{*}$ and $p=d/\binom{n-1}{r-1}$, the graph $H\sim\mathcal{H}_{r}(n,p)$ satisfies
@@
$a_{k}=\frac{|S_{k}|}{\Phi},\quad b_{k}=\frac{|S_{k}\setminus\cup_{l<k}S_{\ell}|}{\Phi},\quad\text{and}\quad c_{k}=\frac{|\cup_{k}S_{k}|}{\Phi}.$
By (H2), (H3), and Theorem 3.8, we have
@@
$b_{k}$ $\in\left[\frac{\varepsilon}{4}\left(\frac{1}{(r-1)}\right)^{\frac{1}{r-1}},\,\frac{\varepsilon}{2}\left(\frac{1}{(r-1)}\right)^{\frac{1}{r-1}}\right],\text{ and }$
$c$ $\leq a_{1}+\frac{\varepsilon(K-1)}{2}\left(\frac{1}{(r-1)}\right)^{\frac{1}{r-1}}$
Thus, $c$ is bounded from above by a constant $C(\varepsilon,K)$ that does not depend on $d$.
@@
($~{}~{}~{}~{}~{}{n\choose a_{1}\Phi}\prod_{k=2}^{K}{n\choose b_{k}\Phi}{c\Phi\choose(a_{k}-b_{k})\Phi}$
$\leq\left(\frac{en}{a_{1}\Phi}\right)^{a_{1}\Phi}\prod_{k=2}^{K}\left(\frac{en}{b_{k}\Phi}\right)^{b_{k}\Phi}\left(\frac{en}{(a_{k}-b_{k})}\right)^{(a_{k}-b_{k})\Phi}$
@@
Now for a fixed $\{S_{k}\}$ satisfying (H2) and (H3), we take union bound over all $(T+1)^{K}$ possible choices of $\{t_{k}\}$ and compute an upper bound on the probability that (H1) is satisfied. Let $E$ be the number of edges $e$ for which there exists $k$ such that all vertices of $e$ lie inside of $S_{k}$. We compute a lower bound on $E$ as follows:
$E$ =($={a_{1}\Phi\choose r}+\sum_{k=2}^{K}{a_{k}\Phi\choose r}-{(a_{k}-b_{k})\Phi\choose r}$
$=\frac{\Phi^{r}}{r!}\left(a_{1}^{r}+\sum_{k=2}^{K}a_{k}^{r}-(a_{k}-b_{k})^{r}+o(1)\right)$
$=\frac{\Phi^{r}}{r!}\left(a_{1}^{r}+\sum_{k=2}^{K}b_{k}\Big{(}(a_{k}-b_{k})^{r-1}+a_{k}(a_{k}-b_{k})^{r-2}+\cdots+a_{k}^{r-1}\Big{)}+o(1)\right)$
$\geq\frac{\Phi^{r}}{r!}\left(a_{1}^{r}+\sum_{k=2}^{K}rb_{k}(a_{k}-b_{k})^{r-1}+o(1)\right)$
where the first term counts edges within $S_{1}$ and the $k$th term counts edges crossing $S_{k}$ and $\cup_{\ell<k}S_{\ell}$. For fixed $\{S_{k}\}$ and $\{t_{k}\}$, (H1) occurs when at least $E$ independent non-edges occur in the sampling of $\{A^{(t)}\}$; this happens with probability at most
@@
$\leq\sup_{\{a_{k}\},\{b_{k}\}}\exp\left(\frac{1}{r-1}\Phi\log d\left(a_{1}+\sum_{k=2}^{K}b_{k}+o_{d}(1)\right)-\frac{1}{r}\Phi\log d\left(a_{1}^{r}+\sum_{k=2}^{K}rb_{k}(a_{k}-b_{k})^{r-1}+o(1)\right)\right)$
$\leq\sup_{\{a_{k}\},\{b_{k}\}}\exp\left(\Phi\log d\left(\frac{a_{1}}{r-1}-\frac{a_{1}^{r}}{r}+\sum_{k=2}^{K}\frac{b_{k}}{r-1}-{b_{k}}(a_{k}-b_{k})^{r-1}+o_{d}(1)+o(1)\right)\right)$
$=\sup_{\{a_{k}\},\{b_{k}\}}\exp\left(\Phi\log d\left(\frac{a_{1}}{r-1}-\frac{a_{1}^{r}}{r}-\sum_{k=2}^{K}{b_{k}}\left((a_{k}-b_{k})^{r-1}-\frac{1}{r-1}\right)+o_{d}(1)+o(1)\right)\right)$
@@
$\sup_{a\in\mathbb{R}}\frac{a}{r-1}-\frac{a^{r}}{r}=\frac{1}{r}\left(\frac{1}{r-1}\right)^{\frac{1}{r-1}}.$
@@
$\leq n^{2K}(T+1)^{K}\exp\left(\Phi\log d\left(-\frac{1}{r(r-1)}\left(\frac{1}{r-1}\right)^{\frac{1}{r-1}}+o_{d}(1)+o(1)\right)\right)$
$=\exp\Big{(}-\Omega(n)\Big{)}$
@@
$\mathbb{E}[N]$ =($={|S|\choose b\Phi}\left(1-\frac{d}{{n-1\choose r-1}}\right)^{b\Phi\choose r}$
@@
$=\exp\left(\Phi\log d\left(-\frac{b^{r}}{r}+o(1)+o_{d}(1)\right)\right)$
$=\exp\big{(}-\Omega(n)\big{)}$
@@ 3.3 Proof of the Intractability Result
Fix $\epsilon>0$, and let $K=\left\lceil\frac{2^{r+1}}{\varepsilon^{r}}\right\rceil+1$, $T=(K-1)m$, $c=\frac{\varepsilon}{32\xi(1+\varepsilon)}$ and $\eta=\frac{\varepsilon}{16}\left(\frac{\log d}{d}\right)^{\frac{1}{r-1}}$. Let $d_{0}=d_{0}(\varepsilon)$ large enough to apply Theorem 3.8, Proposition 3.9, and Lemma 3.10. As usual, we let $\Phi:=\left(\frac{\log d}{d}\right)^{\frac{1}{r-1}}n$.
@@
First we show that $|V_{t}\triangle V_{t-1}|\leq\frac{\varepsilon}{4}\left(\frac{1}{r-1}\cdot\frac{\log d}{d}\right)^{\frac{1}{r-1}}n$. From (E1) we know that the failure event setting $V^{\eta}_{f}=\varnothing$ does not occur for all $t\in[T]$. By Definition 1.1, $i\in V_{t}\triangle V_{t-1}$ if one of two events occur:
@@
$\mathbb{E}_{A\sim\mathcal{H}_{r}(n,p)}\big{[}\|f(A)\|^{2}\big{]}\leq\xi(1+\varepsilon)\Phi$
@@
Combining with Lemma LABEL:Lemma:FMM_upper_bound, we have $(1+\varepsilon)\Phi\leq|V_{t}\cap L|\leq\left(r^{\frac{1}{r-1}}+\varepsilon\right)\Phi$ with probability at least $1-\delta-\exp\{-\Omega(n)\}$. Now suppose that for some $0\leq T^{\prime}\leq T-m$, we have sampled $A^{(0)},\;dots,A^{(T^{\prime})}$ and $S_{1}=V_{0},\cdots,S_{K^{\prime}}=V_{t_{k^{\prime}}}$ have been successfully selected. As mentioned previously, $A^{(T^{\prime}+m)}$ is independent from $A^{(0)},\ldots,A^{(T^{\prime})}$. So, provided $|S_{k}|\leq\left(r^{\frac{1}{r-1}}+\varepsilon\right)\Phi$ for $1\leq k\leq K^{\prime}$, we may apply Lemma 3.10 with $S=\cup_{k\leq K^{\prime}}S_{k}$ and $a=\left(r^{\frac{1}{r-1}}+\varepsilon\right)K^{\prime}$. As a result, $|V_{T^{\prime}+m}\cap(\cup_{k\leq K^{\prime}}S_{k})|\leq\varepsilon\Phi$ with probability $1-\exp\{-\Omega(n)\}$. Therefore, we have
@@
This implies that we can find $S_{K^{\prime}+1}=V_{t}$ for some $t\leq T^{\prime}+m$ and thus by induction the process succeeds by timestep $T=(K-1)m$. By a union bound over $t$, (E1) holds with probability at least $1-\delta(T+1)-\exp\{-\Omega(n)\}$.
By Definition 2.4 and the stability of $f$, (E2) holds with probability at least $\left(d/{n-1\choose r-1}\right)^{4(K-1)D/c}$. By Proposition 3.9, (E3) holds with probability $1-\exp\{-\Omega(n)\}$. We claim that it suffices to have
$\left(\frac{d}{{n-1\choose r-1}}\right)^{4(K-1)D/c}\geq 2\exp\{-Cn\},\quad\text{ and }\quad\left(\frac{d}{{n-1\choose r-1}}\right)^{4(K-1)D/c}\geq 2\delta Kn^{r}$ (7)
@@
$\left(\frac{d}{{n-1\choose r-1}}\right)^{4(K-1)D/c}$ $\geq\exp\{-Cn\}+\delta Kn^{r}$
$>\exp\{-\Omega(n)\}+\delta Km$
$>\exp\{-\Omega(n)\}+\delta(T+1).$
@@
And this is implied by $D<\frac{C_{1}n}{\xi\log n}$, for large enough $n$ and some constant $C_{1}=C_{1}(\varepsilon,d)>0$. The second inequality is implied by $\delta\leq\exp\left(-\frac{4(K-1)(r-1)D}{c}\log n-r\log n-\log(2K)\right)$. For large enough $n$, given $\xi,D\geq 1$, there exists another constant $C_{2}=C_{2}(\varepsilon,d)>0$ such that the prior inequality is implied by $\delta\leq\exp\left(-C_{2}\xi D\log n\right)$. ∎

### v3

@@ 3.1.2 From Local to Low-Degree
Here, $\cong$ denotes the isomorphism of rooted hypergraphs. The following lemma shows that $n_{T}$ is well-concentrated.
@@
The proof follows by combining Markov’s inequality with a hypergraph version of Proposition 12.3 [Bar+19] (for which the proof is identical, mutatis mutandis). ∎
@@
The proof follows by a hypergraph version of [Bar+19, Lemma 12.4] (for which the proof is identical, mutatis mutandis), and linearity of expectation. ∎
@@
For any $\tau>0$ and $n$ sufficiently large in terms of d,s,r,d,\,s,\,r, and $\tau$, there is a constant $C\coloneqq C(d,s,r,\tau)>0$ such that
@@
Before we prove the main result, we make an analogous observation to the one in [Wei22, Lemma 2.11] on graphs. In particular, a random polynomial can be converted to a deterministic polynomial that works almost as well.
@@
Thus, there exists an $\omega^{*}\in\Omega$ such that the deterministic degree-$D$ polynomial $f^{*}(\cdot)=f(\cdot,\omega)$ satisfies
$\mathbb{E}_{A}[\|f^{*}(A)\|^{2}]\leq c\xi k\quad\text{and}\quad\mathbb{P}_{A}[|V_{f}^{\eta}(A)|<k]\leq c\delta$
@@
Given $\varepsilon>0$, let $d$ and $s$ be sufficiently large in order to apply Theorem 3.1. As a result, we obtain a valid $s$-local function $g$ that outputs independent sets of density
@@
which proves that the rounding procedure succeeds with probability at least $1-\exp(-\Omega(n^{1/3}))$.
@@
The idea now is to partition $V_{q}$ into subsets $U_{1},\ldots,U_{(r-1)q+1}$ such that for each $U_{i}$, the vertices in $U_{i}$ have disjoint $s$-neighborhoods. In particular, the random variables $\{\mathbbm{1}_{v\in I_{f}}\}_{v\in U_{i}}$ are independent conditioned on the outcome $A$. Each vertex $v\in V_{q}$ has at most (r1)q(r-1)q vertices in $N_{2s}(A,v)$, which implies there are at most (r1)q(r-1)q vertices $u\in V_{q}$ such that $u\neq v$ and $N_{s}(A,v)\cap N_{s}(A,u)\neq\varnothing$. We can greedily form the partition $\{U_{i}\}$ while maintaining the disjointness property (given that there are (r1)q+1(r-1)q+1 many such subsets available). For such a partition $\{U_{i}\}$, we apply Chernoff’s bound for each $i$ to obtain the following:
@@
It suffices to show that $\mathbb{E}_{A,\mathbf{X}}[f_{v}(A,\mathbf{X})^{2}]=O(1)$ for all vertices $v$. Fix a vertex $v$ and let $N=|N_{s}(A,v)|$. By (3), for each $H\in\mathbb{H}_{v,s,q}$ the corresponding term is not 0 only if $H$ is a subgraph of $N_{s}(A,v)$. This implies that number of of nonzero terms is at most
${N\choose\leq q}=\sum_{i=0}^{q}{N\choose i}\leq\sum_{i=1}^{q}N^{i}\leq(N+1)^{q}.$
@@
In order to bound the $\mathbb{E}_{A,\mathbf{X}}[f_{v}(A,\mathbf{X})^{2}]$, it is sufficient to prove a tail bound for $N$. Starting from $m_{0}=1$, let $m_{i}$ be the number of vertices whose distance from $v$ in $A$ is exactly $i$. Conditioned on $m_{i}$, $m_{i+1}$ is stochastically dominated by $\mathsf{Bin}\left(r{n-1\choose r-1}m_{i},d/{n-1\choose r-1}\right)$. By a Chernoff bound, for fixed $m_{i}\geq 1$ and any $\delta\geq 1$, we have
$\mathbb{P}[m_{i+1}\geq(1+\delta)rdm_{i}]\leq\exp\left(-\frac{\delta rdm_{i}}{3}\right)\leq\exp\left(-\frac{\delta rd}{3}\right).$
Note that this inequality holds trivially if $m_{i}=0$. Taking a union bound over all $i\in[s]$,
@@
For $t=(3\delta rd)^{s}$, we have a subexponentially small tail bound on $N$:
@@ 3.2 Forbidden Structures
In this section, we will define two “forbidden structures” in order to assist with our proof of Theorem 1.6. Before we define these structures and prove the corresponding results, we state the following theorem due to Krivelevich and Sudakov on the size of the maximum independent set in $\mathcal{H}_{r}(n,p)$ (this result wasn’t explicitly stated, but can inferred from Corollary 1 in their paper) [KS98].
@@
Let $\varepsilon>0$ and $r\in\mathbb{N}$ such that $r\geq 2$. There is $d_{0}>0$ such that for any $d\geq d_{0}$, there exists $n_{0}>0$ such that for any $n\geq n_{0}$ and $p=d/\binom{n-1}{r-1}$, the graph $H\sim\mathcal{H}_{r}(n,p)$ satisfies
@@
$a_{k}=\frac{|S_{k}|}{\Phi},\quad b_{k}=\frac{|S_{k}\setminus\cup_{l<k}S_{\ell}|}{\Phi},\quad\text{and}\quad c=\frac{|\cup_{k}S_{k}|}{\Phi}.$
By (H2) and (H3), we have
@@
$b_{k}$ $\in\left[\frac{\varepsilon}{4}\left(\frac{1}{r-1}\right)^{\frac{1}{r-1}},\,\frac{\varepsilon}{2}\left(\frac{1}{r-1}\right)^{\frac{1}{r-1}}\right]$
Given the even event in conclusion of Theorem 3.8 holds with probability $1-e^{-\Omega(n)}$, we have
$c\leq a_{1}+\frac{\varepsilon(K-1)}{2}\left(\frac{1}{(r-1)}\right)^{\frac{1}{r-1}},$
with probability $1-e^{-\Omega(n)}$. In other words, $c$ is bounded from above by a constant $C(\varepsilon,K)$ that does not depend on $d$.
@@
$~~~~~{n\choose a_{1}\Phi}\prod_{k=2}^{K}{n\choose b_{k}\Phi}{c\Phi\choose(a_{k}-b_{k})\Phi}$
$\leq\left(\frac{en}{a_{1}\Phi}\right)^{a_{1}\Phi}\prod_{k=2}^{K}\left(\frac{en}{b_{k}\Phi}\right)^{b_{k}\Phi}\left(\frac{ec}{(a_{k}-b_{k})}\right)^{(a_{k}-b_{k})\Phi}$
@@
Now for a fixed $\{S_{k}\}$ satisfying (H2) and (H3), we take union bound over all $(T+1)^{K}$ possible choices of $\{t_{k}\}$ and compute an upper bound on the probability that (H1) is satisfied. Let $E$ be the number of potential edges $e$ for which there exists $k$ such that all vertices of $e$ lie inside of $S_{k}$. We compute a lower bound on $E$ as follows:
$E$ $\geq{a_{1}\Phi\choose r}+\sum_{k=2}^{K}{a_{k}\Phi\choose r}-{(a_{k}-b_{k})\Phi\choose r}$
$=\frac{\Phi^{r}}{r!}\left(a_{1}^{r}+\sum_{k=2}^{K}a_{k}^{r}-(a_{k}-b_{k})^{r}+o_{d}(1)\right)$
$=\frac{\Phi^{r}}{r!}\left(a_{1}^{r}+\sum_{k=2}^{K}b_{k}\Big((a_{k}-b_{k})^{r-1}+a_{k}(a_{k}-b_{k})^{r-2}+\cdots+a_{k}^{r-1}\Big)+o_{d}(1)\right)$
$\geq\frac{\Phi^{r}}{r!}\left(a_{1}^{r}+\sum_{k=2}^{K}rb_{k}(a_{k}-b_{k})^{r-1}+o_{d}(1)\right)$
In the first line, the first term counts potential edges within $S_{1}$ and the $k$th term counts potential edges crossing $S_{k}$ and $\cup_{\ell<k}S_{\ell}$. Notice that equality holds when all the $t_{k}$ are equal, and otherwise the number of non-edges that need to occur is only larger. For fixed $\{S_{k}\}$ and $\{t_{k}\}$, (H1) occurs when at least $E$ independent non-edges occur in the sampling of $\{A^{(t)}\}$; this happens with probability at most
@@
$\leq\sup_{\{a_{k}\},\{b_{k}\}}\exp\left(\frac{1}{r-1}\Phi\log d\left(a_{1}+\sum_{k=2}^{K}b_{k}+o_{d}(1)\right)-\frac{1}{r}\Phi\log d\left(a_{1}^{r}+\sum_{k=2}^{K}rb_{k}(a_{k}-b_{k})^{r-1}+o_{d}(1)\right)\right)$
$\leq\sup_{\{a_{k}\},\{b_{k}\}}\exp\left(\Phi\log d\left(\frac{a_{1}}{r-1}-\frac{a_{1}^{r}}{r}+\sum_{k=2}^{K}\frac{b_{k}}{r-1}-{b_{k}}(a_{k}-b_{k})^{r-1}+o_{d}(1)\right)\right)$
$=\sup_{\{a_{k}\},\{b_{k}\}}\exp\left(\Phi\log d\left(\frac{a_{1}}{r-1}-\frac{a_{1}^{r}}{r}-\sum_{k=2}^{K}{b_{k}}\left((a_{k}-b_{k})^{r-1}-\frac{1}{r-1}\right)+o_{d}(1)\right)\right)$
@@
$\sup_{a\in\mathbb{R}}\left(\frac{a}{r-1}-\frac{a^{r}}{r}\right)=\frac{1}{r}\left(\frac{1}{r-1}\right)^{\frac{1}{r-1}}.$
@@
$\leq n^{2K}(T+1)^{K}\exp\left(\Phi\log d\left(-\frac{1}{r(r-1)}\left(\frac{1}{r-1}\right)^{\frac{1}{r-1}}+o_{d}(1)\right)\right)$
$=\exp\left(-\Omega(n)\right),$
@@
$\mathbb{E}[N]$ $={|S|\choose b\Phi}\left(1-\frac{d}{{n-1\choose r-1}}\right)^{b\Phi\choose r}$
@@
$=\exp\left(\Phi\log d\left(-\frac{b^{r}}{r}+o_{d}(1)\right)\right)$
$=\exp\left(-\Omega(n)\right),$
@@ 3.3 Proof of the Intractability Result
Fix $\epsilon>0$, and let $K=\left\lceil\frac{2^{r+1}}{\varepsilon^{r}}\right\rceil+1$, T=(K1)mT=(K-1)m, $c=\frac{\varepsilon}{32\xi(1+\varepsilon)}$ and $\eta=\frac{\varepsilon}{16}\left(\frac{\log d}{(r-1)d}\right)^{\frac{1}{r-1}}$. Let $d_{0}=d_{0}(\varepsilon)$ large enough to apply Theorem 3.8, Proposition 3.9, and Lemma 3.10. For the convenience of this proof, we let $\Phi:=\left(\frac{\log d}{(r-1)d}\right)^{\frac{1}{r-1}}n$, which differs from the previous definition only by a multiplicative constant.
@@
First we show that $|V_{t}\triangle V_{t-1}|\leq\frac{\varepsilon}{4}\left(\frac{1}{r-1}\cdot\frac{\log d}{d}\right)^{\frac{1}{r-1}}n=\frac{\varepsilon}{4}\Phi$. From (E1) we know that the failure event setting $V^{\eta}_{f}=\varnothing$ does not occur for all $t\in[T]$. By Definition 1.1, $i\in V_{t}\triangle V_{t-1}$ if one of two events occur:
@@
$\mathbb{E}_{A\sim\mathcal{H}_{r}(n,p)}\big[\|f(A)\|^{2}\big]\leq\xi(1+\varepsilon)\Phi.$
@@
Combining with Theorem 3.8, we have $(1+\varepsilon)\Phi\leq|V_{t}|\leq\left(r^{\frac{1}{r-1}}+\varepsilon\right)\Phi$ with probability at least $1-\delta-\exp(-\Omega(n))$. Now suppose that for some $0\leq T^{\prime}\leq T-m$, we have sampled $A^{(0)},\ldots,A^{(T^{\prime})}$ and $S_{1}=V_{0},\ldots,S_{K^{\prime}}=V_{t_{K^{\prime}}}$ have been successfully selected for some $K^{\prime}<K$ and $t_{K^{\prime}}\leq T^{\prime}$. As mentioned previously, $A^{(T^{\prime}+m)}$ is independent from $A^{(0)},\ldots,A^{(T^{\prime})}$. So provided $|S_{k}|\leq\left(r^{\frac{1}{r-1}}+\varepsilon\right)\Phi$ for $1\leq k\leq K^{\prime}$, we may apply Lemma 3.10 with $S=\cup_{k\leq K^{\prime}}S_{k}$ and $a=\left(r^{\frac{1}{r-1}}+\varepsilon\right)\left(\frac{1}{r-1}\right)^{\frac{1}{r-1}}K^{\prime}$. As a result, $|V_{T^{\prime}+m}\cap(\cup_{k\leq K^{\prime}}S_{k})|\leq\varepsilon\Phi$ with probability $1-\exp(-\Omega(n))$. Therefore, we have
@@
This implies that we can find $S_{K^{\prime}+1}=V_{t}$ for some $t\leq T^{\prime}+m$ and thus by induction the process succeeds by timestep T=(K1)mT=(K-1)m. By a union bound over $t$, (E1) holds with probability at least $1-\delta(T+1)-\exp(-\Omega(n))$.
By Definition 2.5 and the stability of $f$, (E2) holds with probability at least $\left(d/{n-1\choose r-1}\right)^{4(K-1)D/c}$. By Proposition 3.9, (E3) holds with probability $1-\exp(-\Omega(n))$. We claim that it suffices to have
$\left(\frac{d}{{n-1\choose r-1}}\right)^{4(K-1)D/c}\geq 2\exp(-Cn),\quad\text{ and }\quad\left(\frac{d}{{n-1\choose r-1}}\right)^{4(K-1)D/c}\geq 2\delta Kn^{r}$ (7)
@@
$\left(\frac{d}{{n-1\choose r-1}}\right)^{4(K-1)D/c}$ $\geq\exp(-Cn)+\delta Kn^{r}$
$>\exp(-\Omega(n))+\delta Km$
$>\exp(-\Omega(n))+\delta(T+1).$
@@
And this is implied by $D<\frac{C_{1}n}{\xi\log n}$, for large enough $n$ and some constant $C_{1}=C_{1}(\varepsilon,d)>0$. The second inequality is implied by
$\delta\leq\exp\left(-\frac{4(K-1)(r-1)D}{c}\log n-r\log n-\log(2K)\right).$ (8)
For large enough $n$, given $\xi,D\geq 1$, there exists another constant $C_{2}=C_{2}(\varepsilon,d)>0$ such that (8) is implied by $\delta\leq\exp\left(-C_{2}\xi D\log n\right)$. ∎